determine analytically if the following functions are even, odd or neither show work
f(x) = 2x^3 − x
1
Expert's answer
2019-04-05T09:41:37-0400

Function "f" is even if and only if "f(-x) = f(x)" for all "x".

Function "f" is odd if and only if "f(-x) = -f(x)" for all "x".


For given function:

"f(-x) = 2(-x)^3 - (-x) = -2x^3 + x = -(2x^3-x) = -f(x)"


Answer: function is odd.